The solution of the initial value problem dy/dx=-x/y, y(4)=3 is
1
Expert's answer
2019-11-25T12:16:40-0500

Let's solve this equation using separate variables method.

Rearange terms:

"ydy = -xdx\\"

Integrate both sides of equation:

"\dfrac12 y^2 = -\dfrac12x^2+C\\"

Multiply by 2 and rearange terms again:

"x^2 + y^2 = C\\"

Obtain constant C from the initial conditions:

"4^2 + 3^2 = C\\
C = 5^2 = 25\\"

Thus, the final solution is:

"x^2 + y^2 = 25\\"

This is the equation of circle with radius 5.